# arXiv:2603.02656v1[quant-ph] 03 Mar 2026

### 03 Mar 2026 v1

@@ 6.4 Completing the Lower Bound Proof
Case 1: query to $A_{G}$.  Under both $\mathcal{D}_{\text{yes}}$ and $\mathcal{D}_{\text{no}}$, $G\sim\mathcal{G}(n,1/2)$, so previously unqueried entries of $A_{G}$ are i.i.d. $\mathrm{Bernoulli}(1/2)$, independent of $\mathcal{F}_{t-1}$.
Case 2: query $(A_{H})_{ij}$ with no collision.  Under $\mathcal{D}_{\text{no}}$, $(A_{H})_{ij}$ is an independent $\mathrm{Bernoulli}(1/2)$ by construction. Under $\mathcal{D}_{\text{yes}}$, $(A_{H})_{ij}=(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}\oplus B_{ij}$ where $B_{ij}\sim\mathrm{Bernoulli}(\varepsilon/2)$. The “no collision” condition means the entry $(\pi^{-1}(i),\pi^{-1}(j))$ has not been queried in $G$. Since $\pi$ is uniformly random and independent of $G$, this $G$-entry is an unqueried $\mathrm{Bernoulli}(1/2)$ variable, independent of $\mathcal{F}_{t-1}$. XOR-ing with $B_{ij}$ (which is also independent of $\mathcal{F}_{t-1}$) gives $(A_{H})_{ij}\sim\mathrm{Bernoulli}(1/2)$, identically under both distributions. ∎
By Subclaim 6.7.1, the transcripts under $\mathcal{D}_{\text{yes}}$ and $\mathcal{D}_{\text{no}}$ are identically distributed conditioned on no collision occurring. To bound the collision probability, consider the $t$-th query to $H$ at position $(i,j)$. Under $\mathcal{D}_{\text{yes}}$, this probes the $G$-entry at position $(\pi^{-1}(i),\pi^{-1}(j))$. Since $\pi$ is uniformly random and independent of the algorithm’s choices, the pair $(\pi^{-1}(i),\pi^{-1}(j))$ is a uniformly random element of $\binom{[n]}{2}$, independent of which $G$-entries have been previously queried. A collision occurs only if this random position coincides with one of the $\leq q_{G}$ previously queried $G$-entries, which happens with probability $\leq q_{G}/\binom{n}{2}$. By a union bound over all $q_{H}$ queries to $H$:
$\Pr[\text{any collision occurs}]\;\leq\;\sum_{t=1}^{q_{H}}\frac{q_{G}}{\binom{n}{2}}\;\leq\;\frac{q_{G}\cdot q_{H}}{\binom{n}{2}}\;\leq\;\frac{q^{2}}{n^{2}/2}\;=\;\mathcal{O}(q^{2}/n^{2}).$
Note that this matches the bound from Lemma 6.6, which is tight. By Subclaim 6.7.1, the total variation distance is at most the collision probability (since non-colliding queries have identical distributions), giving:
$d_{\mathrm{TV}}(\mathcal{T}_{\text{yes}},\mathcal{T}_{\text{no}})\;\leq\;\mathcal{O}(q^{2}/n^{2}).$
For the advantage to exceed $1/3$, we need $q^{2}/n^{2}=\Omega(1)$, i.e., $q=\Omega(n)$.
[…]
For constant $\varepsilon\in(0,1/8)$, the direct argument gives $q^{2}/n^{2}=\Omega(1)$, hence $q=\Omega(n)$. To reach $\Omega(n^{2})$, we use a birthday-paradox refinement: the algorithm needs $\Omega(\sqrt{\binom{n}{2}})=\Omega(n)$ queries to $H$ just to have a constant probability of any collision. But each collision reveals only $\mathcal{O}(1)$ bits, and the algorithm needs $\Omega(n)$ bits of information about $\pi$ to distinguish the cases (since $\pi$ has $\Omega(n\log n)$ bits of entropy and the algorithm must determine whether $\pi$ is planted or the graphs are independent). This information- theoretic argument gives $q=\Omega(n^{2})$ for $\varepsilon=\Theta(1)$. ∎
@@ 7.1 Spectral Similarity
The algorithm proceeds by quantum eigenvalue estimation. Using the block-encoding framework [GSL+19], we construct a block encoding of $L_{G}/\lambda_{\max}$ from the adjacency oracle $O_{G}$ with $\mathcal{O}(1)$ queries per application.
@@
Remark: For eigenvalues with small multiplicity, the coupon-collector overhead can be avoided using the eigenvalue sampling algorithm of Gilyén et al. [GSL+19], which estimates the empirical spectral distribution directly at cost $\mathcal{O}(\sqrt{n}/\eta)$ per sample. For our purposes, the simpler repeated-QPE approach suffices.
@@ 9.5 Connections to Graph Neural Networks
Recent advances in graph neural networks (GNNs) have highlighted the importance of the Weisfeiler–Leman (WL) hierarchy for graph isomorphism testing [MRF+19, XHL+19]. The $k$-WL test is known to be equivalent in power to the counting logic $\mathcal{C}^{k+1}$ [CFI92]. An interesting direction is whether quantum walks on the product graph can be connected to the WL hierarchy—specifically, whether the information obtained from $t$ steps of the quantum walk corresponds to refinements in a particular level of the WL hierarchy.

### 28 Mar 2026 v2

@@ 6.4 Completing the Lower Bound Proof
Case 1: query to $A_{G}$. Under both $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, $G\sim\mathcal{G}(n,\tfrac{1}{2})$ independently of $\mathcal{F}_{t-1}$ and of $H$. Any previously unqueried entry $(A_{G})_{ij}$ is therefore $\mathrm{Bern}(\tfrac{1}{2})$, independent of $\mathcal{F}_{t-1}$, under both distributions.
Case 2: query $(A_{H})_{ij}$, conditioned on no collision at step $t$. Under $D_{\mathrm{no}}$, $(A_{H})_{ij}\sim\mathrm{Bern}(\tfrac{1}{2})$ independently of everything, so the claim holds immediately.
Under $D_{\mathrm{yes}}$, $(A_{H})_{ij}=(A_{G})_{\pi^{-1}(i)\,\pi^{-1}(j)}\oplus B_{ij}$ where $B_{ij}\sim\mathrm{Bern}(\varepsilon/2)$ is independent of $G$, $\pi$, and $\mathcal{F}_{t-1}$. The no-collision condition means $(\pi^{-1}(i),\pi^{-1}(j))\notin\mathcal{Q}_{G}$, the set of positions of $A_{G}$ already queried. Since $\pi\sim\mathrm{Uniform}(S_{n})$ is independent of $G$, of $\mathcal{Q}_{G}$, and of the algorithm’s history, the random position $(\pi^{-1}(i),\pi^{-1}(j))$ is uniform over $\binom{[n]}{2}$; conditioned on it falling outside $\mathcal{Q}_{G}$ (the no-collision event), it is uniform over the at least $\binom{n}{2}-|\mathcal{Q}_{G}|$ unqueried positions of $A_{G}$. Each such position holds an independent $\mathrm{Bern}(\tfrac{1}{2})$ value (by the $\mathcal{G}(n,\tfrac{1}{2})$ structure), independent of $\mathcal{F}_{t-1}$. XOR-ing with the independent noise bit $B_{ij}\sim\mathrm{Bern}(\varepsilon/2)$:
$(A_{H})_{ij}\;\sim\;\mathrm{Bern}\!\Bigl(\tfrac{1}{2}\bigl(1-\tfrac{\varepsilon}{2}\bigr)+\tfrac{1}{2}\cdot\tfrac{\varepsilon}{2}\Bigr)\;=\;\mathrm{Bern}\!\bigl(\tfrac{1}{2}\bigr),$
which is identical to the $D_{\mathrm{no}}$ distribution. ∎
Part 1: $\Omega(n)$ lower bound (average-case).
By Lemma 6.6 and Subclaim 6.6.1, the total-variation distance between the transcript distributions satisfies
$d_{\mathrm{TV}}(T_{\mathrm{yes}},T_{\mathrm{no}})\;\leq\;\frac{4\varepsilon q^{2}}{n^{2}}.$
By Lemmas 6.3 and 6.5, $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are supported on valid YES and NO instances respectively with probability $1-e^{-\Omega(n^{2})}$. For any algorithm to achieve success probability $\geq 2/3$ on both distributions, its distinguishing advantage must be at least $1/3$. Since the advantage is at most $d_{\mathrm{TV}}(T_{\mathrm{yes}},T_{\mathrm{no}})+e^{-\Omega(n^{2})}$, we need
$\frac{4\varepsilon q^{2}}{n^{2}}\;\geq\;\frac{1}{3}-e^{-\Omega(n^{2})}\;\geq\;\frac{1}{4}\qquad\text{for large }n,$
which gives $q\geq n\,/\,(4\sqrt{3\varepsilon})=\Omega(n)$ for constant $\varepsilon$.
Part 2: $\Omega(n^{2})$ lower bound (worst-case adaptive argument).
We prove that for any constant $\varepsilon\in(0,\tfrac{1}{8})$ and any deterministic adaptive algorithm $\mathcal{A}$ making $q<\varepsilon\binom{n}{2}/2$ queries in total, there exist a YES instance and a NO instance that $\mathcal{A}$ cannot distinguish.
Lemma 6.7 (Hard-instance construction).
Let $G^{*}$ be any graph with $|E(G^{*})|\geq\binom{n}{2}/4$ (which holds for $G^{*}\sim\mathcal{G}(n,\tfrac{1}{2})$ with probability $1-e^{-\Omega(n^{2})}$). For any deterministic adaptive algorithm $\mathcal{A}$ making $q_{H}<\varepsilon\binom{n}{2}/2$ queries to $A_{H}$, there exist graphs $H_{\mathrm{yes}}$ and $H_{\mathrm{no}}$ on $[n]$ such that:
  1. (a)
    $\mathrm{ed}(G^{*},H_{\mathrm{yes}})=0$  (a valid YES instance with $\bar{\mathrm{ed}}=0\leq\varepsilon$);
  2. (b)
    $\mathrm{ed}(G^{*},H_{\mathrm{no}})\geq 2\varepsilon\binom{n}{2}$ (a valid NO instance with $\bar{\mathrm{ed}}\geq 2\varepsilon$);
  3. (c)
    $\mathcal{A}$ produces the same query transcript on $(G^{*},H_{\mathrm{yes}})$ and $(G^{*},H_{\mathrm{no}})$.
Proof.
Since $G^{*}$ is fixed, $\mathcal{A}$’s queries to $A_{G}$ return determined values. $\mathcal{A}$’s queries to $A_{H}$ are adaptive but still form a well-defined sequence of positions once $H$ is specified. We construct $H_{\mathrm{yes}}$, then $H_{\mathrm{no}}$.
Constructing $H_{\mathrm{yes}}$. Set $H_{\mathrm{yes}}=G^{*}$ (the identity isomorphism). Then $\mathrm{ed}(G^{*},H_{\mathrm{yes}})=0$. Running $\mathcal{A}$ on $(G^{*},H_{\mathrm{yes}})$ produces a sequence of $q_{H}$ adaptive queries to $A_{H_{\mathrm{yes}}}$ at positions $\mathcal{S}_{H}=\{(i_{1},j_{1}),\ldots,(i_{q_{H}},j_{q_{H}})\}\subseteq\binom{[n]}{2}$, with answers $\mathbf{a}=(a_{1},\ldots,a_{q_{H}})\in\{0,1\}^{q_{H}}$. (These are well-defined since $H_{\mathrm{yes}}=G^{*}$ is fixed.)
Constructing $H_{\mathrm{no}}$. Define $H_{\mathrm{no}}$ as the graph with adjacency matrix:
$(A_{H_{\mathrm{no}}})_{ij}\;=\;\begin{cases}a_{k}&\text{if }(i,j)=(i_{k},j_{k})\in\mathcal{S}_{H},\\ 0&\text{otherwise.}\end{cases}$
That is, $H_{\mathrm{no}}$ agrees with $H_{\mathrm{yes}}$ on every queried position and has no edges outside $\mathcal{S}_{H}$.
Verifying property (c). Since $H_{\mathrm{no}}$ returns the same answer $a_{k}$ at every position $(i_{k},j_{k})\in\mathcal{S}_{H}$, and these are exactly the positions $\mathcal{A}$ queries adaptively (the adaptive query sequence is the same because each query $(i_{k},j_{k})$ and its answer $a_{k}$ are identical), $\mathcal{A}$ produces the same transcript on both inputs.
Verifying property (b). For any permutation $\pi\in S_{n}$:
$\mathrm{ed}_{\pi}(G^{*},H_{\mathrm{no}})$ $=\bigl|\{(i,j)\in\tbinom{[n]}{2}:(A_{G^{*}})_{ij}\neq(A_{H_{\mathrm{no}}})_{\pi(i)\pi(j)}\}\bigr|$
$\geq\bigl|\{(i,j):(A_{G^{*}})_{ij}=1\text{ and }(\pi(i),\pi(j))\notin\mathcal{S}_{H}\}\bigr|$
$\geq|E(G^{*})|-|\mathcal{S}_{H}|$
$\geq\tfrac{1}{4}\tbinom{n}{2}-q_{H}\;\geq\;\tfrac{1}{4}\tbinom{n}{2}-\tfrac{\varepsilon}{2}\tbinom{n}{2}\;=\;\bigl(\tfrac{1}{4}-\tfrac{\varepsilon}{2}\bigr)\tbinom{n}{2}$.
For $\varepsilon<\tfrac{1}{8}$, this is at least $\tfrac{1}{8}\binom{n}{2}>2\varepsilon\binom{n}{2}$. Since the bound holds for every $\pi$, we have $\mathrm{ed}(G^{*},H_{\mathrm{no}})=\min_{\pi}\mathrm{ed}_{\pi}(G^{*},H_{\mathrm{no}})\geq 2\varepsilon\binom{n}{2}$. ∎
Completing the $\Omega(n^{2})$ proof.
By Lemma 6.7, for every deterministic algorithm $\mathcal{A}$ with $q<\varepsilon\binom{n}{2}/2$ queries, and for a typical $G^{*}$ (satisfying $|E(G^{*})|\geq\binom{n}{2}/4$, which holds with probability $1-e^{-\Omega(n^{2})}$ for $G^{*}\sim\mathcal{G}(n,\tfrac{1}{2})$), there exist a YES instance $(G^{*},H_{\mathrm{yes}})$ and a NO instance $(G^{*},H_{\mathrm{no}})$ on which $\mathcal{A}$ produces the same transcript. Therefore $\mathcal{A}$ outputs the same answer for both, and must err on at least one.
Since every deterministic algorithm with $q<\varepsilon\binom{n}{2}/2$ queries fails on at least one valid input, by Yao’s minimax principle [Yao77] any randomized algorithm with $q<\varepsilon\binom{n}{2}/2$ queries fails with probability at least $\tfrac{1}{2}>\tfrac{1}{3}$ on at least one of $\{(G^{*},H_{\mathrm{yes}}),(G^{*},H_{\mathrm{no}})\}$.
Since $\varepsilon\binom{n}{2}/2=\Omega(n^{2})$ for constant $\varepsilon$, any randomized algorithm achieving success probability $\geq\tfrac{2}{3}$ on all valid inputs requires $q=\Omega(n^{2})$ queries. ∎
Remark 6.8.
The Part 1 and Part 2 arguments are complementary. Part 1 shows that the planted correlated model $D_{\mathrm{yes}}$ vs. $D_{\mathrm{no}}$ is hard on average, with the difficulty scaling as $\Omega(n/\sqrt{\varepsilon})$. Part 2 shows that specific worst-case instances require $\Omega(n^{2})$ queries regardless of the algorithm’s strategy: a graph with $\Omega(n^{2})$ edges cannot be distinguished from an empty-looking adversary-chosen graph without reading a constant fraction of $A_{H}$. Together they establish Theorem 6.1 in full. The quantum upper bound of $O(n^{3/2}\log n/\varepsilon)$ (Theorem 5.9) therefore yields a polynomial quantum speedup of $\tilde{\Omega}(\sqrt{n})$ in the worst-case query model.
@@ 7.1 Spectral Similarity
The algorithm proceeds by quantum eigenvalue estimation. Using the block-encoding framework [GSLW19], we construct a block encoding of $L_{G}/\lambda_{\max}$ from the adjacency oracle $O_{G}$ with $\mathcal{O}(1)$ queries per application.
@@
Remark: For eigenvalues with small multiplicity, the coupon-collector overhead can be avoided using the eigenvalue sampling algorithm of Gilyén et al. [GSLW19], which estimates the empirical spectral distribution directly at cost $\mathcal{O}(\sqrt{n}/\eta)$ per sample. For our purposes, the simpler repeated-QPE approach suffices.
@@ 9.5 Connections to Graph Neural Networks
Recent advances in graph neural networks (GNNs) have highlighted the importance of the Weisfeiler–Leman (WL) hierarchy for graph isomorphism testing [MRF+19, XHLJ19]. The $k$-WL test is known to be equivalent in power to the counting logic $\mathcal{C}^{k+1}$ [CFI92]. An interesting direction is whether quantum walks on the product graph can be connected to the WL hierarchy—specifically, whether the information obtained from $t$ steps of the quantum walk corresponds to refinements in a particular level of the WL hierarchy.
@@ 9.6 Broader Impact
Acknowledgments
The author thanks Clément Canonne for his careful reading and insightful feedback, which substantially improved both the presentation and the substance of this work—particularly by clarifying the importance of explicitly formulating the correlated-input model and by prompting a more complete and rigorous treatment of the classical lower-bound proof.